I need to find out the number of foxes and rabbits. Thanks
Foxes and rabbits. The ratio of foxes to rabbits in the Deerfield Forest Preserve is 2 to 9. If there are 35 fewer foxes than rabbits, then how many of each are there?
Let x=rabbits
Let x-35=foxes

Answer by scott8148(6628)

(x - 35) / x = 2 / 9

"cross" multiplying ___ 9x - 315 = 2x

7x = 315 ___ x = 45